# JOGS: Joint Optimization of Pose Estimation and 3D Gaussian Splatting

### v1

@@ Abstract
Traditional novel view synthesis methods heavily rely on external camera pose estimation tools such as COLMAP, which often introduce computational bottlenecks and propagate errors. To address these challenges, we propose a unified framework that jointly optimizes 3D Gaussian points and camera poses without requiring pre-calibrated inputs. Our approach iteratively refines 3D Gaussian parameters and updates camera poses through a novel co-optimization strategy, ensuring simultaneous improvements in scene reconstruction fidelity and pose accuracy. The key innovation lies in decoupling the joint optimization into two interleaved phases: first, updating 3D Gaussian parameters via differentiable rendering with fixed poses, and second, refining camera poses using a customized 3D optical flow algorithm that incorporates geometric and photometric constraints. This formulation progressively reduces projection errors, particularly in challenging scenarios with large viewpoint variations and sparse feature distributions, where traditional methods struggle. Extensive evaluations on multiple datasets demonstrate that our approach significantly outperforms existing COLMAP-free techniques in reconstruction quality, and also surpasses the standard COLMAP-based baseline in general.
1 Introduction
@@
Recent advancements in the field of computer vision have led to significant progress in 3D scene reconstruction and rendering. In particular, the introduction of 3D Gaussian Splatting (3DGS) [1] technology has provided an efficient and realistic technique for scene representation and rendering. 3DGS explicitly models the scene using a group of Gaussian ellipsoids. This provides rapid and accurate rendering, clearly exhibiting its benefits in real-time situations. Due to its explicit representation and efficient rendering capabilities, 3DGS has been widely applied in various fields, especially in scenarios requiring efficient processing and realistic rendering [2, 3].
@@
Accurate pose estimation is extremely important [4] for most novel view synthesis methods including 3DGS. Most existing 3DGS methods do not include a pose estimation component, but rely on external inputs (e.g., COLMAP [5, 6]). The separation of pose estimation and 3DGS optimization may lead to suboptimal solutions. On the other hand, the reliance on external input may limit its application to certain scenarios [7, 4]. To solve these issues, recent research suggests many 3DGS solutions that do not require inputs of camera poses. For example, CFGS [4] uses a combined optimization of camera parameters and Gaussians. This method transforms the camera pose registration problem into an image optimization task between two consecutive frames. It achieves excellent reconstruction results in continuous and dense image streams, but meet problems when the constraints are violated. ZeroGS [8] and InstantSplat [9] do not require pre-supplied camera poses, but they need to load a pre-trained model in advance for pose estimation, and usually work in very sparse views. These methods either require the integration of additional information or pre-trained models, or can only operate with strong constraints images, making the application scenarios are very limited.
@@
In this paper, to address these challenges, we propose a unified framework that jointly optimizes the camera poses and the 3D Gaussian representations without requiring pre-calibrated inputs. In addition to the losses used in the standard 3DGS algorithm [1], our framework introduces a reprojection loss to penalize the inconsistencies between different views. After an initial coarse pipeline setup, the camera poses are subsequently optimized in conjunction with the 3DGS parameters using the alternating direction method (ADM) algorithm [10]. In each iteration, the 3DGS parameters are updated following the standard 3DGS algorithm. For the refinement of camera poses, we propose a Lucas-Kanade 3D optical flow (LK3D) algorithm, which leverages Gaussians and image reprojection errors by integrating image gradients with transformation-based projection error relationships. This alternating optimization strategy significantly improves pose accuracy and achieves stable convergence even under large camera viewpoint movements or sparse feature distributions.
@@
For validation, we compare the proposed method with several state-of-the-art methods on three public datasets, including Tanks and Temples [11], LLFF-NeRF [12] and Shiny [13]. The experimental results show that our method outperforms compared methods in novel view synthesis, and achieves high reconstruction quality in different scenarios. The source code of our method will be released upon paper publication.
@@
(2) we propose an LK3D algorithm to optimize camera poses based on the reprojection errors between 3D Gaussians and image pixels, which is independent of the sequential relationship between images and is able to effectively fine-tune the camera pose; and
@@ 2 Related Work
Figure 1: Method Overview. Our JOGS framework jointly optimizes Pose Estimation and 3D Gaussian Splatting. It starts with a simple SfM initialization, then iteratively updates 3D Gaussian splatting parameters $\mathcal{G}$ and refines camera poses $\mathcal{P}$, ensuring simultaneous improvements in scene reconstruction fidelity and pose accuracy. The updating of Gaussian points follows a standard 3DGS pipeline, while the refinement of camera poses is done by the proposed LK3D algorithm.
@@
Novel View Synthesis(NVS). The task aims to generate photorealistic renderings of target scenes from unknown viewpoints using a limited set of input images. Recent advancements in neural rendering have significantly improved NVS in terms of reconstruction quality and efficiency. The seminal work on Neural Radiance Fields (NeRF) [14] introduced a paradigm shift in NVS by representing scenes as continuous implicit neural radiance fields, encoded via multilayer perceptrons (MLPs). Subsequent studies have extended NeRF in various directions:  Barron et al. [15, 16] focused on fundamental enhancements. Some works [17, 18, 19, 20, 21, 22] enhanced dynamic scene modeling. Some methods [23, 24] optimized computational efficiency to accelerate training, and some works [25, 26] integrated AIGC with NeRF to facilitate few-shot or zero-shot 3D scene generation. However, NeRF continues to face challenges, including prolonged training times, high hardware demands and limited editability. Recently, the emergence of 3D Gaussian Splatting (3DGS) [1] has achieved breakthroughs by utilizing explicit differentiable representations, striking a balance between rendering quality and efficiency. Extensive research has been conducted on 3DGS, covering areas such as scene rendering quality and realism [27, 28, 29], 3DGS acceleration [30, 31], geometry reconstruction [32, 33], dynamic scenes [34, 35, 36] and few-shot reconstruction [37, 38, 39]. Nevertheless, most existing methods still rely on camera poses and sparse point clouds precomputed by COLMAP [6, 5].
@@
NVS without Pose Input. Eliminating the dependence of input pose has become a main topic in recent research of NVS, for both NeRF and 3DGS methods. I-NeRF [40] introduced inverse rendering to estimate camera poses through keypoint alignment using pre-trained NeRF. BARF [41] proposed a coarse-to-fine coordinate encoding strategy, with further improvement in GARF [42, 43]. Nope-NeRF [7] trained NeRF by incorporating undistorted depth priors. For 3DGS-based methods, CFGS [4] is the most closely related to our work. It builds the entire 3D Gaussian in a continuous fashion, "growing" some Gaussian points with each new view added. It optimizes the camera pose by minimizing the photometric loss between the rendered image and the next frame image. While it achieves 3DGS scene representation without relying on COLMAP, its optimization depends on the temporal relationship between adjacent images, and the change of view angles between consecutive frames needs to be small. ZeroGS [8] relies on a pre-trained DUSt3R-based [44, 45] model called Spann3R [46]. InstantSplat [9] implements camera-free pose reconstruction in sparse views. While GSHT [47] achieves quality enhancement over CFGS, this improvement is constrained by the inherent reliance of the method on the temporal ordering within image sequences. In summary, current mainstream methods either require the integration of additional information or pre-trained models [40, 7, 8, 44, 45], hence limited to working with only a small number of images due to high computational resource [41, 7, 9], or assume minimal camera motion [7, 4, 47]. To overcome these limitations, we design a new framework that jointly optimizes 3D Gaussian and camera pose.
3 Method
3.1 Problem Definition
@@
In traditional 3D Gaussian splatting methods, $\mathcal{P}$ is treated as known parameters, and the problem is reduced to:
@@
In this paper, we treat both $\mathcal{G}$ and $\mathcal{P}$ as learnable parameters, and optimize them jointly in the training step. To this end, we introduce a 3D optical flow loss to penalize the difference between the projections of two different views. The definition and the optimization method are described in detail in Section˜3.4.
3.2 Joint Optimization Framework
@@
Our joint optimization framework establishes a dual-phase alternating minimization scheme to solve the coupled problem in Section˜3.1. Let $\mathcal{G}^{(t)}$ and $\mathcal{P}^{(t)}=\{R_{i}^{(t)},\textbf{s}_{i}^{(t)}\}_{i=1}^{n}$ denote the 3D Gaussian parameters and camera poses at iteration $t$. As shown in Algorithm˜1, these two parts of parameters are optimized by an alternating direction method, which contains two phases as follows:
@@
Specifically, the incremental pose adjustment is computed using the LK3D algorithm described in Algorithm˜2, of which the detailed explanation is described in Section˜3.4.
@@
3.3 Initialize Camera Poses and Gaussian Points
@@
Unlike the standard 3DGS, our method reconstructs both the initial sparse point cloud and camera poses entirely from scratch, without relying on external tools or pose priors. This design ensures compatibility with our joint optimization pipeline and provides greater control over reconstruction quality, sparsity, and initialization behavior.
3.4 Camera Pose Refinement
We propose a method of optimizing camera pose based on 3D Gaussians and image reprojection error. As illustrated in Algorithm˜1, our method interleaves camera pose refinement with 3D Gaussian Splatting (3DGS) training during the initial $m$ iterations. Specifically, we freeze the Gaussian parameters when performing pose optimization at regular intervals, while updating the 3DGS parameters in the remaining iterations. This alternating strategy ensures stable gradient propagation for subsequent scene reconstruction. During the training process, the camera pose was optimized with the 3DGS model, and the camera rotation matrix and shift vector were calculated using the projection of a 3D Gaussian points from multiple viewpoints.
@@ 3.4 Camera Pose Refinement
Lucas-Kanade 3D Optical Flow Algorithm. Let $P=[R|\mathbf{s}]$ denote the camera pose of a certain target image. Given a Gaussian point $g\in\mathcal{G}$, we denote $c(g)$ the color value of $g$ and $\mathbf{x}(g)=(x_{g},y_{g},z_{g})^{\top}$ the 3D position coordinates of $g$ in the world coordinate system. We then define a transformation function $\mathcal{W}(\mathbf{x}(g);P)$ that maps the 3D Gaussian coordinates $\mathbf{x}(g)$ from the world coordinate system to the target image plane following the standard projective geometry.
@@
$P^{*}=\arg\min_{P}\sum_{g\in\mathcal{G}}\big(c(g)-I(\mathcal{W}(\mathbf{x}(g);P))\big)^{2}.$ (6)
By minimizing the differences in pixels between the source and transformed images, the optimal pose parameters $P^{*}$ can be determined.
It is difficult to directly compute the optimal camera pose $P$, since no close-form solution is available. Our method uses a gradient-based update approach by extending the standard LK algorithm [52] to 3-dimensional space, which iteratively revises the transformation matrix $P$ with an increment $\Delta P$ as:
$\Delta P^{*}=\arg\min_{\Delta P}\sum_{g\in\mathcal{G}}\big(I(\mathcal{W}(\mathbf{x}(g);P+\Delta P))-c(g)\big)^{2}.$ (7)
@@
$\Delta P^{*}\approx\arg\min_{\Delta P}\sum_{g\in\mathcal{G}}\big(I(\mathcal{W}(\mathbf{x}(g);P))+\nabla I\frac{\partial\mathcal{W}}{\partial P}\Delta P-c(g)\big)^{2},$ (8)
where $\nabla I$ represents the image gradient and $\frac{\partial\mathcal{W}}{\partial P}$ is the Jacobian matrix of the transformation function $\mathcal{W}$ with respect to the transformation matrix $P$. According to the principle that the derivative at the extreme value is zero, the pose increment $\Delta P$ is computed via Gauss-Newton approximation:
$\Delta P\approx H^{-1}\sum_{g\in\mathcal{G}}\Big(\nabla I\frac{\partial\mathcal{W}}{\partial P}\Big)^{\top}\big(c(g)-I(\mathcal{W}(\mathbf{x}(g);P))\big),$ (9)
@@
$H=\sum_{g\in\mathcal{G}}\left(\nabla I\frac{\partial\mathcal{W}}{\partial P}\right)^{\top}\left(\nabla I\frac{\partial\mathcal{W}}{\partial P}\right).$ (10)
@@
4 Experiments
4.1 Experimental Setup
Datasets. We conducted extensive experiments on various datasets, including LLFF-NeRF [12], Tanks and Temples [11] and Shiny [13]. LLFF-NeRF: This dataset contains real-world multi-view images captured by various devices, comprising eight scenes. The number of images varies across scenes, with the scene fern having the fewest (twenty images) and horns the most (sixty-two images). Tanks and Temples: This dataset comprises eight scenes, encompassing both indoor and outdoor environments. In line with the configurations of CFGS [4] and Nope-NeRF [7], we further enhanced the complexity of the dataset. Given the limited variation in camera poses in the original dataset, we uniformly sampled one-fifth of the images from each scene to amplify the pose variation between consecutive frames. Shiny: The dataset consists of a number of challenging scenes with significant reflected or refracted lighting changes. Since the data volume per scene varies across datasets, we adopted proportional data splitting rather than using fixed quantities. For each scene, seven-eighths of the data were allocated for training, with the remaining one-eighth reserved for testing.
Evaluation Metrics. We employed the same evaluation metrics as CFGS and Nope-NeRF. For novel view synthesis, we used standard metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [53], and Learned Perceptual Image Patch Similarity (LPIPS) [54]. For pose evaluation, we treated COLMAP-estimated poses as ground truth and measured Absolute Trajectory Error (ATE), which includes Relative Rotation Error (RPEr) and Relative Translation Error (RPEt), along with Relative Pose Error (RPE). ATE quantifies the discrepancy between estimated camera positions and ground truth, while RPE measures relative pose errors between image pairs.
Implementation Details. We initialized camera poses and sparse Gaussian points using only scene images and camera intrinsics. During 3D Gaussian reconstruction, we alternately optimized 3D Gaussians and camera poses. Global pose optimization was performed every 100 Gaussian iterations, limited to the first 15,000 iterations. This restricted optimization strategy prevents error accumulation, as pose estimation errors could degrade reconstruction quality, which might further corrupt pose estimation accuracy. Thus, optimizing poses only during the initial quarter of training iterations is empirically justified. All experiments were conducted on a single RTX 3090 GPU. Unless otherwise stated, our experiments follow the same 3DGS parameter settings.
4.2 Comparing with Baseline
Our experimental framework is built upon the original 3D Gaussian Splatting (3DGS) architecture [1]. While the proposed modules are theoretically compatible with advanced 3DGS variants, our current implementation specifically adheres to the canonical formulation due to two methodological considerations: (1) make sure that the comparison with the original 3DGS can be made directly so that future generations can easily reproduce our work; and (2) isolating the performance impact of our contributions from other confounding factors. To maintain consistency, all architectural parameters strictly follow the original 3DGS configuration. This design choice facilitates direct comparability with COLMAP-based 3DGS baselines under identical experimental protocols.
Table 1: Quantitative comparison on Tanks and Temples. The best results are highlighted in bold, and the second in underline, and the same styles are adopted in the subsequent tables.
@@ 4.2 Comparing with Baseline
Table 2: Quantitative comparison on LLFF-NeRF. For the Fortress and Leaves scenes (marked with *), we directly cite the results of CFGS and GSHT from zeroGS [8], because our experimental environment could not meet the running requirements of their codes.
@@
Table 3: Quantitative comparison on Shiny. The original dataset contains eight scenes, but both CFGS and GSHT suffer from running errors and fail to report the final results in some scenes. Thus we only report the experimental results for four scenes.
Scene PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
3DGS CFGS GSHT Ours 3DGS CFGS GSHT Ours 3DGS CFGS GSHT Ours
Cd 28.29 26.60 26.44 28.18 0.94 0.87 0.90 0.94 0.12 0.17 0.16 0.12
Giants 21.69 14.37 16.01 20.52 0.72 0.45 0.36 0.68 0.28 0.47 0.62 0.25
Lab 28.54 26.26 27.99 29.28 0.93 0.82 0.91 0.94 0.15 0.18 0.15 0.15
Tools 24.59 12.44 11.67 24.33 0.84 0.50 0.47 0.83 0.35 0.59 0.60 0.32
Mean 25.77 18.51 19.05 25.58 0.86 0.58 0.59 0.85 0.23 0.39 0.43 0.21

### v2

@@ Abstract
Traditional novel view synthesis methods heavily rely on external camera pose estimation tools such as COLMAP, which often introduce computational bottlenecks and propagate errors. To address these challenges, we propose a unified framework that jointly optimizes 3D Gaussian points and camera poses without requiring pre-calibrated inputs. Our approach iteratively refines 3D Gaussian parameters and updates camera poses through a novel co-optimization strategy, ensuring simultaneous improvements in scene reconstruction fidelity and pose estimation accuracy. The key innovation lies in decoupling the joint optimization into two interleaved phases: first, updating 3D Gaussian parameters via differentiable rendering with fixed poses, and second, refining camera poses using a customized 3D optical flow algorithm that incorporates geometric and photometric constraints. This formulation progressively reduces projection errors, particularly in challenging scenarios with large viewpoint variations and sparse feature distributions, where traditional methods struggle. Extensive evaluations on multiple datasets demonstrate that our approach significantly outperforms existing COLMAP-free techniques in reconstruction quality, and also surpasses the standard COLMAP-based baseline in general.
I Introduction
@@
Recent advancements in the field of computer vision have led to significant progress in 3D scene reconstruction and rendering. In particular, the introduction of 3D Gaussian Splatting (3DGS) [1] technology has provided an efficient and realistic technique for scene representation and rendering. 3DGS explicitly models the scene using a group of Gaussian ellipsoids. This provides rapid and accurate rendering, clearly exhibiting its benefits in real-time situations. Due to its explicit representation and efficient rendering capabilities, 3DGS has been widely applied in various fields [2, 3, 4, 5, 6, 7], especially in scenarios requiring efficient processing and realistic rendering [8].
@@
Accurate pose estimation is extremely important [9] for most novel view synthesis methods including 3DGS. Most existing 3DGS methods do not include a pose estimation component, but rely on external inputs (e.g., COLMAP [10, 11]). The separation of pose estimation and 3DGS optimization may lead to suboptimal solutions. On the other hand, the reliance on external input may limit its application to certain scenarios [12, 9]. To solve these issues, recent research suggests many 3DGS solutions that do not require inputs of camera poses. For example, CFGS [9] uses a combined optimization of camera parameters and Gaussian points. This method transforms the camera pose registration problem into an image optimization task between two consecutive frames. It achieves excellent reconstruction results in continuous and dense image streams, but its performance degradations when the constraints are violated. ZeroGS [13] and InstantSplat [14] do not require pre-supplied camera poses, but they need to load a pre-trained model in advance for pose estimation, and usually work in very sparse views. These methods either require the integration of additional information or pre-trained models, or can only operate with strong constraints on the input images, which substantially limits their applicability.
@@
In this paper, to address these challenges, we propose JOGS that Jointly Optimizes the camera poses and the 3D Gaussian Splatting representations without requiring pre-calibrated inputs. In addition to the losses used in the standard 3DGS algorithm [1], our framework introduces a reprojection loss to penalize the inconsistencies between different views. After an initial coarse pipeline setup, the camera poses are subsequently optimized in conjunction with the 3DGS parameters using the alternating direction method (ADM) algorithm [15]. In each iteration, the 3DGS parameters are updated following the standard 3DGS algorithm. For the refinement of camera poses, we propose a lucas-kanade 3D optical flow (LK3D) algorithm, which leverages Gaussian points and image reprojection errors by integrating image gradients with transformation-based projection error relationships. This alternating optimization strategy significantly improves pose accuracy and achieves stable convergence even under large camera viewpoint movements or sparse feature distributions.
@@
For validation, we compare the proposed method with several state-of-the-art methods on three public datasets, including Tanks and Temples [16], LLFF-NeRF [17] and Shiny [18]. The experimental results show that our method outperforms the baselines in novel view synthesis, and achieves high reconstruction quality in different scenarios. The source code of our method will be released upon paper publication.
@@
(2) we propose an LK3D algorithm to optimize camera poses based on the reprojection errors between 3D Gaussian points and image pixels, which is independent of the sequential relationship between images and is able to effectively fine-tune the camera pose; and
@@ II Related Work
Figure 1: Method Overview. Our JOGS framework jointly optimizes Pose Estimation and 3D Gaussian Splatting. It starts with a simple SfM initialization, then iteratively updates 3D Gaussian splatting parameters $\mathcal{G}$ and refines camera poses $\mathcal{P}$, ensuring simultaneous improvements in scene reconstruction fidelity and pose estimation accuracy. The updating of Gaussian points follows a standard 3DGS pipeline, while the refinement of camera poses is done by the proposed LK3D algorithm.
II-A Novel View Synthesis
@@
The task aims to generate photorealistic renderings of target scenes from unknown viewpoints using a limited set of input images. Recent advancements in neural rendering [19, 20, 21, 22] have significantly improved NVS in terms of reconstruction quality and efficiency. The seminal work on Neural Radiance Fields (NeRF) [19] introduced a paradigm shift in NVS by representing scenes as continuous implicit neural radiance fields, encoded via multilayer perceptrons (MLPs). Subsequent studies extended NeRF along several directions. For example, Mip-NeRF 360 [23] improve core rendering fidelity and scene representation. Some works [24, 25] enhanced dynamic scene modeling. Some methods [26, 27] optimized computational efficiency to accelerate training. However, NeRF continues to face challenges, including prolonged training times, high hardware demands and limited editability. Recently, the emergence of 3DGS [1] has achieved breakthroughs by utilizing explicit differentiable representations, striking a balance between rendering quality and efficiency. Extensive research has been conducted on 3DGS, covering areas such as scene rendering quality and realism [28, 29], 3DGS acceleration [30], geometry reconstruction [31, 32], dynamic scenes [33] and few-shot reconstruction [34, 35]. Nevertheless, most existing methods still rely on camera poses and sparse point clouds precomputed by COLMAP [11, 10].
II-B NVS without Pose Input
@@
Eliminating the dependence of input pose has become a main topic in recent research of NVS, for both NeRF and 3DGS methods. I-NeRF [36] introduced inverse rendering to estimate camera poses through keypoint alignment using pre-trained NeRF. BARF [37] proposed a coarse-to-fine coordinate encoding strategy, with further improvement in GARF [38, 39]. Nope-NeRF [12] trained NeRF by incorporating undistorted depth priors. For 3DGS-based methods, CFGS [9] is the most closely related to our work. It builds the entire 3D Gaussian in a continuous fashion, ”growing” some Gaussian points with each new view added. It optimizes the camera pose by minimizing the photometric loss between the rendered image and the next frame image. While it achieves 3DGS scene representation without relying on COLMAP, its optimization depends on the temporal relationship between adjacent images, and the change of view angles between consecutive frames needs to be small. ZeroGS [13] relies on a pre-trained DUSt3R-based [40] model called Spann3R [41]. InstantSplat [14] implements camera-free pose reconstruction in sparse views. While GSHT [42] achieves quality enhancement over CFGS, this improvement is constrained by the inherent reliance of the method on the temporal ordering within image sequences. In summary, current mainstream methods either require the integration of additional information or pre-trained models [13, 40], hence limited to working with only a small number of images due to high computational resource [37, 12, 14], or assume minimal camera motion [12, 9, 42]. To overcome these limitations, we design a new framework that jointly optimizes 3D Gaussian and camera pose.
III Method
III-A Problem Definition
@@
In traditional 3DGS methods, $\mathcal{P}$ is treated as known parameters, and the problem is reduced to:
@@
In this paper, we treat both $\mathcal{G}$ and $\mathcal{P}$ as learnable parameters, and optimize them jointly in the training step. To this end, we introduce a 3D optical flow loss to penalize the difference between the projections of two different views. The definition and the optimization method are described in detail in Section III-D.
III-B Joint Optimization Framework
@@
Our joint optimization framework establishes a dual-phase alternating minimization scheme to solve the coupled problem in Section III-A. Let $\mathcal{G}^{(t)}$ and $\mathcal{P}^{(t)}=\{R_{i}^{(t)},\textbf{s}_{i}^{(t)}\}_{i=1}^{n}$ denote the 3D Gaussian parameters and camera poses at iteration $t$. As shown in Algorithms 1, these two parts of parameters are optimized by an alternating direction method, which contains two phases as follows:
@@
Specifically, the incremental pose adjustment is computed using the LK3D algorithm described in Algorithms 2, of which the detailed explanation is described in Section III-D.
@@
III-C Initialize Camera Poses and Gaussian Points
@@
Unlike the standard 3DGS, our method reconstructs both the initial sparse Gaussian points and camera poses entirely from scratch, without relying on external tools or pose priors. This design ensures compatibility with our joint optimization pipeline and provides greater control over reconstruction quality, sparsity, and initialization behavior.
III-D Camera Pose Refinement
We propose a method of optimizing camera pose based on 3D Gaussian points and image reprojection error. As illustrated in Algorithms 1, our method interleaves camera pose refinement with 3DGS training during the initial $m$ iterations. Specifically, we freeze the Gaussian parameters when performing pose optimization at regular intervals, while updating the 3DGS parameters in the remaining iterations. This alternating strategy ensures stable gradient propagation for subsequent scene reconstruction. During the training process, the camera pose was optimized with the 3DGS model, and the camera rotation matrix and shift vector were calculated using the projection of a 3D Gaussian points from multiple viewpoints.
@@ III-D Camera Pose Refinement
Lucas-Kanade 3D Optical Flow Algorithm. Let $\mathbf{p}=[R|\mathbf{s}]\in\mathcal{P}$ denote the camera pose of a certain target image. Given a Gaussian point $g\in\mathcal{G}$, we denote $\mathbf{c}(g)$ the color value of $g$ and $\mathbf{x}(g)=(x_{g},y_{g},z_{g})^{\top}$ the 3D position coordinates of $g$ in the world coordinate system. We then define a transformation function $\mathcal{W}(\mathbf{x}(g);\mathbf{p})$ that maps the 3D Gaussian coordinates $\mathbf{x}(g)$ from the world coordinate system to the target image plane following the standard projective geometry.
@@
$\mathbf{p}^{*}=\arg\min_{\mathbf{p}}\sum_{g\in\mathcal{G}}\big(\mathbf{c}(g)-I(\mathcal{W}(\mathbf{x}(g);\mathbf{p}))\big)^{2}.$ (6)
By minimizing the differences in pixels between the source and transformed images, the optimal pose parameters $\mathbf{p}^{*}$ can be determined.
It is difficult to directly compute the optimal camera pose $P$, since no close-form solution is available. Our method uses a gradient-based update approach by extending the standard LK algorithm [47] to 3-dimensional space, which iteratively revises the transformation matrix $\mathbf{p}$ with an increment $\Delta\mathbf{p}$ as:
$\Delta\mathbf{p}^{*}=\arg\min_{\Delta\mathbf{p}}\sum_{g\in\mathcal{G}}\big(I(\mathcal{W}(\mathbf{x}(g);\mathbf{p}+\Delta\mathbf{p}))-\mathbf{c}(g)\big)^{2}.$ (7)
@@
$\Delta\mathbf{p}^{*}\approx\arg\min_{\Delta P}\sum_{g\in\mathcal{G}}\big(I(\mathcal{W}(\mathbf{x}(g);\mathbf{p}))+\nabla I\frac{\partial\mathcal{W}}{\partial\mathbf{p}}\Delta\mathbf{p}-\mathbf{c}(g)\big)^{2},$ (8)
where $\nabla I$ represents the image gradient and $\frac{\partial\mathcal{W}}{\partial\mathbf{p}}$ is the Jacobian matrix of the transformation function $\mathcal{W}$ with respect to the transformation parameters $\mathbf{p}$. According to the principle that the derivative at the extreme value is zero, the pose increment $\Delta\mathbf{p}$ is computed via Gauss-Newton approximation:
$\Delta\mathbf{p}\approx H^{-1}\sum_{g\in\mathcal{G}}\Big(\nabla I\frac{\partial\mathcal{W}}{\partial\mathbf{p}}\Big)^{\top}\big(\mathbf{c}(g)-I(\mathcal{W}(\mathbf{x}(g);\mathbf{p}))\big),$ (9)
@@
$H=\sum_{g\in\mathcal{G}}\left(\nabla I\frac{\partial\mathcal{W}}{\partial\mathbf{p}}\right)^{\top}\left(\nabla I\frac{\partial\mathcal{W}}{\partial\mathbf{p}}\right).$ (10)
@@
TABLE I: Quantitative comparison on Tanks and Temples. The best results are highlighted in bold, and the second in underline, and the same styles are adopted in the subsequent tables.
@@
TABLE II: Quantitative comparison on LLFF-NeRF. For the Fortress and Leaves scenes (marked with *), we directly cite the results of CFGS and GSHT from zeroGS [13], because our experimental environment could not meet the running requirements of their codes.
@@
III-E Optimization Strategy of Rotation Matrix.
Specifically, the rotation matrix $R$ is parameterized as:
$R=R_{z}(\gamma)R_{y}(\beta)R_{x}(\alpha),$ (11)
IV Experiments
IV-A Datasets
We conducted extensive experiments on various datasets, including LLFF-NeRF [17], Tanks and Temples [16] and Shiny [18]. LLFF-NeRF: This dataset contains real-world multi-view images captured by various devices, comprising eight scenes. The number of images varies across scenes, with the scene fern having the fewest (twenty images) and horns the most (sixty-two images). Tanks and Temples: This dataset comprises eight scenes, encompassing both indoor and outdoor environments. In line with the configurations of CFGS [9] and GSHT [42], we further enhanced the complexity of the dataset. Given the limited variation in camera poses in the original dataset, we uniformly sampled one-fifth of the images from each scene to amplify the pose variation between consecutive frames. Shiny: The dataset consists of a number of challenging scenes with significant reflected or refracted lighting changes. Since the data volume per scene varies across datasets, we adopted proportional data splitting rather than using fixed quantities. For each scene, seven-eighths of the data were allocated for training, with the remaining one-eighth reserved for testing.